the greater of two numbers is one more than twice the smaller. three times the greater exceeds five times the smaller by ten. find the two numbers.
Answer by jorel1380(3719)

y=2x+1
3y-10=5x
6x+6-10=5x
6x-4=5x
x=4
y=9..